solve equation, check solution
- 4/9 y + 2 = -6

Found 3 solutions by hyperasad, imadewfoos, mathwiz07:
Answer by hyperasad(6)

-4/9 y + 2 = -6
-4/9 y + 2 -2 = -6 -2
-4/9 y = -8
-4/9 y x 9 = -8 x 9
-4 y = -72
-4 y/-4 = -72/-4
y = 18
CHECK
Put the value of y =18 in the equation
-4/9 x 18 + 2 = -6
-72/9 + 2 = -6
-8 + 2 = -6
-6 = -6
LHS = RHS

Answer by imadewfoos(2)

-4/9y+2=-6
____ -2 -2
2 cancels out and you get:
-4/9y=-8
Now you divide:
-8 divided by -4/9=
-8/1 divided by 9/-4 (because you use the reciprocal -4/9=9/-4)
The end product is 18.

Answer by mathwiz07(3)

-4/9y + 2+=-6